OpenStudy (anonymous):

Solve the system of equations for x or y, whichever you do not eliminate 5x + 4y=10 5x -4y=10 a.x=4 b.x=5 c.x=10 d.x=2

OpenStudy (istim):

So, rearrange one of them so that it equals x or y. (So that it becoems either x=#-#y or y=#/#x) Then, sub it into the other equation as x or y (The same variable as the first equation)

OpenStudy (phi):

add these equations together 5x + 4y=10 5x -4y=10 10x + 0=20 solve for x

OpenStudy (anonymous):

Thanks=)

OpenStudy (earthcitizen):

\[\left[\begin{matrix}5 & 4 \\ 5 & -4\end{matrix}\right]\left(\begin{matrix}x \\y\end{matrix}\right)=\left(\begin{matrix}10 \\ 10\end{matrix}\right)\]